# Sirui Xu1  Hung Yu Ling2  Yu-Xiong Wang1†  Liang-Yan Gui1†

### v1

@@ 4 Experiments
Datasets. We use the following datasets: OMOMO [70], BEHAVE [5], HODome [180], IMHD [191], and HIMO [93]. OMOMO, containing 15 objects and approximately 10 hours of data, is our primary dataset for evaluating the full teacher-student distillation framework for its scale. We train 17 teacher policies, one per subject, with subject 14 reserved as the test set and the remaining data used for training the student policy. A small portion of data is discarded after teacher imitation due to severe MoCap errors that could not be corrected (see Sec. F and Sec. H of the supplementary). Additional datasets are used to evaluate teacher policies in various MoCap scenarios with different error levels and interaction types. We focus on highly dynamic motions (Figure 1) and interactions involving multiple body parts (Figure 4), while excluding scenarios such as carrying a bag with a strap, since the simulator [96] used lacks full support for soft bodies.
Metrics. We use the following metrics: (i) Success Rate is defined as the proportion of references that the policy successfully imitates, averaged over all references, while (ii) Duration is the time (in seconds) that the imitation is maintained without triggering the interaction early termination conditions introduced in Sec. 3.2. (iii) Human Tracking Error ($E_{h}$), which measures the per-joint position error (cm) between the simulated and reference human (excluding hand joints for BEHAVE [5] and OMOMO [70] due to inaccuracy), and (iv) the Object Tracking Error ($E_{o}$), which measures the per-point position error (cm) between the simulated and reference object. Both errors are averaged over the duration of the imitation.
@@ 4.1 Quantitative Evaluation
PPO Reference Distillation Policy Distillation Architecture OMOMO-Train [70] OMOMO [70]-Test OMOMO [70]-Test (w ×$\times 10$) HOI-Diff [103] InterDiff [160]
@@ 4.2 Qualitative Evaluation
Figure 4 shows a representative sequence from the experiment in Table 1, illustrating how our teacher policy corrects interactions that PhysHOI [142] fails to track robustly – our method effectively withstands and corrects incorrect hand positioning and floating contacts in the reference. Beyond obvious errors, our method also rectifies the rotation of symmetric objects that MoCap inaccurately depicts as sliding along the ground, shown in Figure 5. Figure 6 presents additional examples that complement Figure 1, demonstrating how our approach integrates with kinematic generators for future interaction prediction and text-to-interaction synthesis. This zero-shot generalization extends to novel objects unseen during training (Figure 7), highlighting the effectiveness of our object geometry and contact-encoded representation, as well as the large-scale training.
@@ 5 Conclusion
{strip}[Uncaptioned image]
Figure A: InterMimic enables simulated humans to perform physical interactions, featuring scalable skill learning covering diverse objects.
@@
    Demo Video. A demonstration video (with a screenshot in Figure A) is provided at demo.mp4, as described in Sec. A.
@@ B Setup of Physical Interaction Simulation
The reference data represent humans using the SMPL models [111, 102]. For simulation, we convert these models into box and cylindrical rigid bodies following [174, 89]. Objects are also converted into simulation models through convex decomposition, as illustrated in Figure B. We summarize the physics parameters for our task in Table A. We follow the physics parameters for the human as specified in [142, 143], with the exception of the specialized range of motion (RoM) for hands, detailed in Table B. Our range of motion (RoM) setting is biologically inspired: finger flexion and extension (bending and straightening) are fully activated. However, unlike the real human, the abduction and adduction of the Metacarpophalangeal (MCP) joint are constrained to minimize the risk of finger interpenetration, in the absence of the correct reference hand pose for guidance. The rationale for these RoM settings is discussed in Sec.3.2 of the main paper and Sec.D.3 of the supplementary.
@@
MCP $[-55.625^{\circ},55.625^{\circ}]$ $[-5.625^{\circ},5.625^{\circ}]$
PIP $[-55.625^{\circ},55.625^{\circ}]$
DIP $[-5.625^{\circ},90.000^{\circ}]$
CMC $[-55.625^{\circ},55.625^{\circ}]$ $[-55.625^{\circ},55.625^{\circ}]$
MCP $[-5.625^{\circ},5.625^{\circ}]$ $[-5.625^{\circ},5.625^{\circ}]$
IP $[-5.625^{\circ},90.000^{\circ}]$ $[-5.625^{\circ},5.625^{\circ}]$
@@ G Additional Experiemental Results
Failure Cases. In Figure F, we illustrate an example where our teacher policies fail to perform successful imitation. Despite the strong adaptability of our policies, as demonstrated in Figures 1 and 5, where they effectively correct reference errors, there are limitations when encountering too many errors. Since the reward design inherently prioritizes reference tracking, excessive errors in the reference inevitably result in failures.
HOI Retargeting. Figure G shows that our teacher policies, trained on reference data for a specific body shape, can successfully drive a human model with a body shape different from the reference in the simulator to accomplish the same task, albeit with slightly varied trajectories. This result highlights the effectiveness of our design, which integrates retargeting into interaction imitation.

### v2

@@ 4 Experiments
Datasets. We use the following datasets from InterAct [101]: OMOMO [40], BEHAVE [3], HODome [109], IMHD [114], and HIMO [53]. OMOMO, containing 15 objects and approximately 10 hours of data, is our primary dataset for evaluating the full teacher-student distillation framework for its scale. We train 17 teacher policies, one per subject, with subject 14 reserved as the test set and the remaining data used for training the student policy. A small portion of data is discarded after teacher imitation due to severe MoCap errors that could not be corrected (see Sec. F and Sec. H of the supplementary). Additional datasets are used to evaluate teacher policies in various MoCap scenarios with different error levels and interaction types. We focus on highly dynamic motions (Figure LABEL:fig:teaser) and interactions involving multiple body parts (Figure 4), while excluding scenarios such as carrying a bag with a strap, since the simulator [54] used lacks full support for soft bodies.
Metrics. We use the following metrics: (i) Success Rate is defined as the proportion of references that the policy successfully imitates at least once, averaged across all references, while (ii) Duration is the time (in seconds) that the imitation is maintained without triggering the interaction early termination conditions introduced in Sec. 3.2. (iii) Human Tracking Error ($E_{h}$) measures the per-joint position error (cm) between the simulated and reference human (excluding hand joints for BEHAVE [3] and OMOMO [40] due to inaccuracy). (iv) Object Tracking Error ($E_{o}$) measures the per-point position error (cm) between the simulated and reference object. Both errors are averaged over the duration of the imitation in the best-performing trial.
@@ 4.1 Quantitative Evaluation
PPO Reference Distillation Policy Distillation Architecture OMOMO-Train [40] OMOMO [40]-Test OMOMO [40]-Test (w $\times 10$) HOI-Diff [62] InterDiff [102]
@@ 4.2 Qualitative Evaluation
Figure 4 shows a representative sequence from the experiment in Table 1, illustrating how our teacher policy corrects interactions that PhysHOI [88] fails to track robustly – our method effectively withstands and corrects incorrect hand positioning and floating contacts in the reference. Beyond obvious errors, our method also rectifies the rotation of symmetric objects that MoCap inaccurately depicts as sliding along the ground, shown in Figure 5. Figure 6 presents additional examples that complement Figure LABEL:fig:teaser, demonstrating how our approach integrates with kinematic generators for future interaction prediction and text-to-interaction synthesis. This zero-shot generalization extends to novel objects unseen during training (Figure 7), highlighting the effectiveness of our object geometry and contact-encoded representation, as well as the large-scale training.
@@ 5 Conclusion
Acknowledgments.
We thank Wei Yang, Yu-Wei Chao, Arsalan Mousavian, Ankur Handa, Samuel Schulter, Morteza Ziyadi, and Xiaohan Fei for valuable discussions. This work was supported in part by NSF Grant 2106825, NIFA Award 2020-67021-32799, the Amazon-Illinois Center on AI for Interactive Conversational Experiences, the Toyota Research Institute, the IBM-Illinois Discovery Accelerator Institute, and Snap Inc. This work used computational resources, including the NCSA Delta and DeltaAI and the PTI Jetstream2 supercomputers through allocations CIS230012, CIS230013, and CIS240311 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, as well as the TACC Frontera supercomputer and Amazon Web Services (AWS) through the National Artificial Intelligence Research Resource (NAIRR) Pilot.
@@ Acknowledgments.
    Demo Video. A demonstration video (with a screenshot in Figure LABEL:fig:teaser_supp) is provided at demo.mp4, as described in Sec. A.
@@ B Setup of Physical Interaction Simulation
The reference data represent humans using the SMPL models [68, 61]. For simulation, we convert these models into box and cylindrical rigid bodies following [51]. Objects are also converted into simulation models through convex decomposition, as illustrated in Figure B. We summarize the physics parameters for our task in Table A. We follow the physics parameters for the human as specified in [88, 89], with the exception of the specialized range of motion (RoM) for hands, detailed in Table B. Our range of motion (RoM) setting is biologically inspired: finger flexion and extension (bending and straightening) are fully activated. However, unlike the real human, the abduction and adduction of the Metacarpophalangeal (MCP) joint are constrained to minimize the risk of finger interpenetration, in the absence of the correct reference hand pose for guidance. The rationale for these RoM settings is discussed in Sec.3.2 of the main paper and Sec.D.3 of the supplementary.
@@
MCP $5^{\circ},55.625^{\circ}] [5.625,5.625$$^{\circ},5.625^{\circ}$]
PIP $55.625^{\circ},55.625^{\circ}]$
DIP $5.625^{\circ},90.000^{\circ}]$
CMC $^{\circ},55.625^{\circ}] [55.625,55.625$$^{\circ},55.625^{\circ}$]
MCP $^{\circ},5.625^{\circ}] [5.625,5.625$$^{\circ},5.625^{\circ}$]
IP $^{\circ},90.000^{\circ}] [5.625,5.625$$^{\circ},5.625^{\circ}$]
@@ G Additional Experiemental Results
Failure Cases. In Figure F, we illustrate an example where our teacher policies fail to perform successful imitation. Despite the strong adaptability of our policies, as demonstrated in Figures LABEL:fig:teaser and 5, where they effectively correct reference errors, there are limitations when encountering too many errors. Since the reward design inherently prioritizes reference tracking, excessive errors in the reference inevitably result in failures.
HOI Retargeting. Figure G shows that teacher policies, trained on reference data for a specific body shape, can successfully drive a human model with a body shape different from the reference in the simulator to accomplish the same task, albeit with slightly varied trajectories. This result highlights the effectiveness of our design, which integrates retargeting into interaction imitation.